# Self-Training: A Survey

### v5

@@ Abstract
Semi-supervised algorithms aim to learn prediction functions from a small set of labeled training set and a large set of unlabeled observations. Because these approaches are relevant in many applications, they have received a lot of interest in both academia and industry. Among the existing techniques, self-training methods have undoubtedly attracted greater attention in recent years. These models are designed to find the decision boundary on low density regions without making additional assumptions about the data distribution, and use the unsigned output score of a learned classifier, or its margin, as an indicator of confidence. The working principle of self-training algorithms is to learn a classifier iteratively by assigning pseudo-labels to the set of unlabeled training samples with a margin greater than a certain threshold. The pseudo-labeled examples are then used to enrich the labeled training data and to train a new classifier in conjunction with the labeled training set. In this paper, we present self-training methods for binary and multi-class classification as well as their variants and two related approaches, namely consistency-based approaches and transductive learning. We also provide brief descriptions of self-supervised learning and reinforced self-training, two distinct approaches despite their similar names. Finally, we present the most popular applications where self-training is employed. For pseudo-labeling, fixed thresholds usually lead to subpar results, highlighting the significance of dynamic thresholding for best results. Moreover, improving pseudo-label noise enhances generalization and class differentiation. The performance is also impacted by augmenting initial labeled training samples. To the best of our knowledge, this is the first thorough and complete survey on self-training.
@@ 1 Introduction
Semi-supervised learning has risen to prominence within the machine learning domain, tackling the core challenge of making inference from both the structure of the unlabeled data and the label information from existing labeled training sets [3]. The framework is particularly useful in scenarios where there are limited labeled examples but an abundance of unlabeled data available for training. This is highly relevant in a range of applications, such as computer vision, natural language processing and speech recognition, where the acquisition of labeled data can be a costly endeavor [103, 21, 35, 74].
@@ 1.2 Three main semi-supervised learning families
Data clustering uses a mixture model and assigns class labels to groups using the labeled data they include; and it constitutes the working principle of generative approaches [66, 49]. The cluster assumption, which underpins these approaches, asserts that if two examples are in the same group, they are likely to be in the same class (Figure 1 (a)). This hypothesis may be explained as follows: if a group is formed by a large number of instances, it is rare that they belong to different classes. This does not imply that a class is constituted by a single group of examples, but rather that two examples from distinct classes are unlikely to be found in the same cluster [1].
If we consider the partitions of instances to be high density areas, a form of the cluster assumption known as low density separation entails determining the decision boundary over low density regions (Figure 1 (b)), and it constitutes the basis of discriminant techniques. The main difference between generative and discriminant techniques is that discriminant approaches find directly the prediction function without making any assumption on how the data are generated [4, 37, 69].
Density estimation is often based on a notion of distance, which may become meaningless for high dimensional spaces. A third hypothesis, known as the manifold assumption, stipulates that in high-dimensional spaces, instances reside on low-dimensional topological spaces that are locally Euclidean (Figure 1 (c)), which is supported by a variety of semi-supervised models called graphical approaches [10, 22].
@@ 1.3 Compatibility
Although semi-supervised algorithms have been successfully applied in many situations, there have been cases where unlabeled data have been shown to have no effect on the performance of a learning task [87]. Several attempts have been made in recent years to investigate the value of unlabeled data in the training process [15, 56], and the capacity of semi-supervised learning approaches to generalize [77, 60]. The bulk of these studies are founded on the notion of compatibility defined by [8], and they strive to exhibit the connection between the marginal data distribution and the target function to be learned. According to these findings, unlabeled data will be beneficial for training only if such a relationship exists.
In generative approaches, the marginal distribution is viewed as a mixture of class conditional distributions, and when compared to the supervised case, semi-supervised learning has been shown to achieve lower finite-sample error bounds in some general cases, or a faster rate of error convergence in others [15, 77, 60, 87]. In this line, [11] showed that accessing the marginal distribution on unlabeled training data would not provide sample size guarantees superior to those obtained by supervised learning unless very strong assumptions about conditional distribution on class labels are made.
For graph-based approaches, [67, 2] provided a context in which such algorithms may be studied and perhaps justified; the key finding of the study is that unlabeled data can help learning in some situations by explicitly defining the structure of the data through a manifold.
@@ 2.1 Semi-supervised framework and notations
In the binary case, $\mathcal{Y}=\{-1,+1\}$, we define the unsigned margin of a classification function $f\in\mathcal{F}$ over an example $\mathbf{x}\in\mathcal{X}$ [23, 5] as
@@
In the multi-class classification case, $\mathcal{Y}=\{1,\ldots,K\}$, the unsigned margin [23, 30] is defined as
@@ 2.2 Self-training: the idea
Self-training, also known as decision-directed or self-taught learning machine, is one of the earliest approach in semi-supervised learning [81, 31] that has risen in popularity in recent years.
To determine the decision boundary on low density regions, the idea behind self-training algorithms is to consider a pseudo-labeling strategy for assigning pseudo-labels to the examples of $X_{\mathcal{U}}$. This strategy can be characterized by a function, called pseudo-labeler:
@@
The self-learning strategy is an iterative wrapper algorithm that starts by learning a supervised classifier on the labeled training set $S$. Then, at each iteration, the current classifier selects a part of the unlabeled data, $X_{\bar{\mathcal{U}}}$, and assigns pseudo-labels to them using the classifier’s predictions.
@@ 2.3 Pseudo-labeling strategies
At each iteration, the self-training selects just a portion of unlabeled data for pseudo-labeling, otherwise, all unlabeled examples would be pseudo-labeled after the first iteration, which would actually result in a classifier with performance identical to the initial classifier [17]. Thus, the implementation of self-training arises the following question: how to determine the subset of examples to pseudo-label?
A classical assumption, that stems from the low density separation hypothesis, is to suppose that the classifier learned at each step makes the majority of its mistakes on observations close to the decision boundary. In the case of binary classification, preliminary research suggested to assign pseudo-labels only to unlabeled observations for which the current classifier is the most confident [91]. Hence, considering thresholds $\theta^{-}$ and $\theta^{+}$ defined for respectively the negative and the positive classes, the pseudo-labeler assigns a pseudo-label $\tilde{y}$ to an instance $\mathbf{x}\in X_{\mathcal{U}}$ such that:
@@
[106] also proposed an adaptation of curriculum learning to pseudo-labeling, which entails in learning a model using easy-to-learn observations before moving on to more complex ones. The principle is that at the step $k$ of the algorithm, the pseudo-labeler selects unlabeled examples having predictions that are in the $(1-\alpha_{k})^{th}$ percentile of the distribution of the maximum probability predictions assumed to follow a Pareto distribution, and where $\alpha_{k}\in[0,1]$ is an hyperparameter that varies from 0 to 1 in increments of $0.2$.
Considering the distribution of predictions over unlabeled data, and the majority vote classifiers, such as Random Forest or Adaboost [79], it is possible to automatically choose a threshold for pseudo-labeling. Formally, the learning of a majority vote classifier with partially labeled data can be defined as follows.
@@
where the denominator is the proportion of the unlabeled examples with the confidence higher than the threshold $\theta$, and the numerator is the joint Bayes risk on $X_{\mathcal{U}}$. Thus, the criterion can be interpreted as a trade-off between the number of examples going to be pseudo-labeled and the error they induce. Furthermore, these bounds are shown to be tight in the case where the majority vote classifier makes its error mostly on low margin regions. [30] demonstrated that this technique outperforms conventional fixed-threshold pseudo-labeling strategies on different multi-class classification problems.
@@
2.4 Self-training with two classifiers
In co-training, examples are defined by two modalities that are comparable but not entirely correlated. Each view of an example is expected to contain complementary information about the data and if there are enough labeled training data, each of them is supposed to be sufficient for learning. The main principle is to learn a classifier on each view, taking initially the available labeled examples as the training set. Then, one of the classifiers assigns pseudo-labels to unlabeled data, and the other one uses them to retrain the model by including them into its training set. At each iteration, the classifiers alternately switch their roles, thereby co-training each other. As for self-training algorithms with a single classifier, this procedure continues until there are no more unlabeled instances to be pseudo-labeled. In practice, several studies artificially generated the two modalities for classification problems where examples are mono-viewed and described by a vector representation. These approaches create the two modalities out of one by selecting at random the set of features that should correspond to each modality from the initial set of features; and their efficiency was empirically proved on various applications [102]. Co-training can thus be thought of as a form of self-training: rather than training a classifier on its own previous predictions, each classifier is trained on the predictions of another classifier that was trained on the predictions of the former. Without splitting the input feature set, [19] proposed Cross Pseudo Supervision for semantic segmentation in images. This method employs two neural networks as supervised classifiers having the same images as inputs. Each neural-network is learned at every mini-batch by considering the pseudo-labels generated by the other network for unlabeled instances as ground-truths. In multi-task learning, [36] proposed to independently train specialized teachers using labeled datasets. These teachers then label an unlabeled dataset, creating a multitask pseudo-labeled dataset. Subsequently, a student model is trained using the pseudo-labeled dataset, employing multi-task learning to learn from various datasets and tasks simultaneously. Finally, the feature representations of the student model are evaluated across six vision tasks, including image recognition, to assess its performance and generalization capabilities.
The learnability of co-training was studied under the PAC framework [92], which also accounts for noise in the class labels of unlabeled examples caused by pseudo-labeling. The injection of noisy labels in the pseudo-labeling step is in fact inherent to any self-training algorithm. Taking into account noisy labels in training a model was first considered in supervised learning, following the paradigm of learning with an imperfect supervisor in which training data contains an unknown portion of imperfect labels [64, 41]. Most of these studies tackle this problem from an algorithmic point of view, employing regularization [62] or estimating mislabeling errors by modeling the transition probability between noisy and true labels [71].
@@ 2.5 Self-training under Domain Shift
When intermediate domains are not given, it is important to ensure that pseudo-labeled target examples are reliable and are not biased towards the source data. While [44] and [110] approached this issue by carefully choosing a pseudo-labeling policy, [78] learn a representation via a tri-training scheme, in which the student is trained on target data pseudo-labeled by agreement of two teachers. [58] alternate between two gradient steps: (1) to train a source classification head that generates pseudo-labels, (2) to train a target classification head using pseudo-labeled data under the constraint that it predicts well on source data.
As the discrepancy between the source and the target can be large, the prediction confidence may exhibit a strong bias failing to distinguish between correct and wrong pseudo-labels. Therefore, several works focus specifically on model calibration and uncertainty estimation including the negative entropy regularization [111], the Monte-Carlo dropout [63], and prediction agreement of diversified linear heads [68].
@@ 3.1 Transductive learning
Transductive approaches also assume that the function class’s structure should reflect prior knowledge of the learning problem at hand, and that it should be built in such a way that the correct prediction of class labels of labeled and unlabeled training examples is contained in a function class $\mathcal{F}_{j}$ of small size with a high probability. In particular, [24] show that with high probability the error on the unlabeled training set of a function from a class function $\mathcal{F}_{j}$ is bounded by its empirical loss on the labeled training set plus a complexity term that depends on the number of labeled examples $m$, the number of unlabeled examples $u$, and the size of the class function $\mathcal{F}_{j}$.
The Transductive Support Vector Machines (TSVM) [45] developed for the binary case is based on this paradigm, and is looking for the optimal hyperplane in a feature space that separates the best labeled examples while avoiding high density areas. TSVM does this by building a structure on a function class $\mathcal{F}$ and sorting the outputs of unlabeled samples by their margins. The solutions to the associated optimization problem are the pseudo-labels of unlabeled examples for which the resulting hyperplane separates the examples of both labeled and unlabeled training sets with the largest margin.

### v6

@@ Abstract
Self-training methods have gained significant attention in recent years due to their effectiveness in leveraging small labeled datasets and large unlabeled observations for prediction tasks. These models identify decision boundaries in low-density regions without additional assumptions about data distribution, using the confidence scores of a learned classifier. The core principle of self-training involves iteratively assigning pseudo-labels to unlabeled samples with confidence scores above a certain threshold, enriching the labeled dataset and retraining the classifier. This paper presents self-training methods for binary and multi-class classification, along with variants and related approaches such as consistency-based methods and transductive learning. We also briefly describe self-supervised learning and reinforced self-training. Furthermore, we highlight popular applications of self-training and discuss the importance of dynamic thresholding and reducing pseudo-label noise for performance improvement.
To the best of our knowledge, this is the first thorough and complete survey on self-training.
@@ 1 Introduction
Self-training is a branch of semi-supervised learning and has emerged as a prominent approach within the machine learning domain, addressing the core challenge of leveraging both labeled and unlabeled data for improved inference. This approach is particularly valuable in scenarios where labeled examples are scarce but there is an abundance of unlabeled data available for training. Self-training has proven highly relevant in a range of applications, such as computer vision, natural language processing, and speech recognition, where the acquisition of labeled data can be costly and time-consuming [93, 19, 29, 64]. By iteratively assigning pseudo-labels to unlabeled samples and retraining the classifier, self-training methods enhance the model’s performance and generalization capabilities.
@@ 1.2 Three main semi-supervised learning families
Data clustering uses a mixture model and assigns class labels to groups using the labeled data they include; and it constitutes the working principle of generative approaches [40]. The cluster assumption, which underpins these approaches, asserts that if two examples are in the same group, they are likely to be in the same class (Figure 1 (a)). This hypothesis may be explained as follows: if a group is formed by a large number of instances, it is rare that they belong to different classes. This does not imply that a class is constituted by a single group of examples, but rather that two examples from distinct classes are unlikely to be found in the same cluster.
If we consider the partitions of instances to be high density areas, a form of the cluster assumption known as low density separation entails determining the decision boundary over low density regions (Figure 1 (b)), and it constitutes the basis of discriminant techniques. The main difference between generative and discriminant techniques is that discriminant approaches find directly the prediction function without making any assumption on how the data are generated [1].
Density estimation is often based on a notion of distance, which may become meaningless for high dimensional spaces. A third hypothesis, known as the manifold assumption, stipulates that in high-dimensional spaces, instances reside on low-dimensional topological spaces that are locally Euclidean (Figure 1 (c)), which is supported by a variety of semi-supervised models called graphical approaches [6].
@@ 1.3 Compatibility
Although semi-supervised algorithms have been successfully applied in many situations, there have been cases where unlabeled data have been shown to have no effect on the performance of a learning task [75]. Several attempts have been made in recent years to investigate the value of unlabeled data in the training process [12], and the capacity of semi-supervised learning approaches to generalize [52]. The bulk of these studies are founded on the notion of compatibility defined by [4], and they strive to exhibit the connection between the marginal data distribution and the target function to be learned. According to these findings, unlabeled data will be beneficial for training only if such a relationship exists.
In generative approaches, the marginal distribution is viewed as a mixture of class conditional distributions, and when compared to the supervised case, semi-supervised learning has been shown to achieve lower finite-sample error bounds in some general cases, or a faster rate of error convergence in others [12, 52, 75]. In this line, [7] showed that accessing the marginal distribution on unlabeled training data would not provide sample size guarantees superior to those obtained by supervised learning unless very strong assumptions about conditional distribution on class labels are made.
For graph-based approaches, [57] provided a context in which such algorithms may be studied and perhaps justified; the key finding of the study is that unlabeled data can help learning in some situations by explicitly defining the structure of the data through a manifold.
@@ 2.1 Semi-supervised framework and notations
In the binary case, $\mathcal{Y}=\{-1,+1\}$, we define the unsigned margin of a classification function $f\in\mathcal{F}$ over an example $\mathbf{x}\in\mathcal{X}$ [2] as
@@
In the multi-class classification case, $\mathcal{Y}=\{1,\ldots,K\}$, the unsigned margin [24] is defined as
@@ 2.2 Self-training: the idea
Self-training, also known as decision-directed or self-taught learning machine, is one of the earliest approach in semi-supervised learning [70] that has risen in popularity in recent years.
To determine the decision boundary on low density regions, the idea behind self-training algorithms is to consider a pseudo-labeling strategy for assigning pseudo-labels to the examples of $X_{\mathcal{U}}$. This strategy can be characterized by a function, called pseudo-labeler [87, 25]:
@@
The self-learning strategy is an iterative wrapper algorithm that starts by learning a supervised classifier on the labeled training set $S$. Then, at each iteration, the current classifier selects a part of the unlabeled data, $X_{\bar{\mathcal{U}}}$, and assigns pseudo-labels to them using the classifier’s predictions [95].
@@ 2.3 Pseudo-labeling strategies
Pseudo-labeling is a crucial component of self-training methods, where a portion of unlabeled data is selected for pseudo-labeling to avoid overfitting to the initial classifier. In the following we will review different techniques that have been proposed to determine the subset of examples to pseudo-label, each with its strengths and weaknesses.
2.3.1 Threshold-Based Methods
A classical assumption, that stems from the low density separation hypothesis, is to suppose that the classifier learned at each step makes the majority of its mistakes on observations close to the decision boundary.
In the case of binary classification, preliminary research suggested to assign pseudo-labels only to unlabeled observations for which the current classifier is the most confident [81]. Hence, considering thresholds $\theta^{-}$ and $\theta^{+}$ defined for respectively the negative and the positive classes, the pseudo-labeler assigns a pseudo-label $\tilde{y}$ to an instance $\mathbf{x}\in X_{\mathcal{U}}$ such that:
@@ 2.3.1 Threshold-Based Methods
These methods are simple and intuitive and they are effective in leveraging high-confidence predictions. However, they present a risk of assigning wrong labels if thresholds are not set appropriately and excessive trust in the confidence measure can be biased due to small labeled sample size.
2.3.2 Proportion-Based Methods
@@
These methods are adaptive to the confidence level of the model and reduce the risk of overfitting to low-confidence predictions. However, they may include incorrect labels if the initial proportion is too high and generally require careful tuning of the proportion parameter.
2.3.3 Curriculum Learning-Based Methods
These methods use curriculum learning to pseudo-label easy-to-learn observations before moving on to more complex ones. [95] proposed an adaptation of curriculum learning to pseudo-labeling, which entails in learning a model using easy-to-learn observations before moving on to more complex ones. The principle is that at the step $k$ of the algorithm, the pseudo-labeler selects unlabeled examples having predictions that are in the $(1-\alpha_{k})^{th}$ percentile of the distribution of the maximum probability predictions assumed to follow a Pareto distribution, and where $\alpha_{k}\in[0,1]$ is an hyperparameter that varies from 0 to 1 in increments of $0.2$. Later on, [21] proposed a method that combines CLIP’s text-image alignment and SAM’s mask generation capabilities, along with a multi-source curriculum learning strategy to address noise and excessive focus issues, gradually improving semantic alignment and segmentation precision.
These approaches gradually increase the complexity of pseudo-labeled data and reduce the risk of incorrect labels in early iterations. However they require careful tuning of the hyperparameter $\alpha_{k}$ and may not be effective if the distribution of predictions is not well-behaved.
2.3.4 Majority vote classifiers
Considering the distribution of predictions over unlabeled data, and the majority vote classifiers, such as Random Forest or Adaboost [68], it is possible to automatically choose a threshold for pseudo-labeling.
Formally, the learning of a majority vote classifier with partially labeled data can be defined as follows.
@@ 2.3.4 Majority vote classifiers
where the denominator is the proportion of the unlabeled examples with the confidence higher than the threshold $\theta$, and the numerator is the joint Bayes risk on $X_{\mathcal{U}}$. Thus, the criterion can be interpreted as a trade-off between the number of examples going to be pseudo-labeled and the error they induce. Furthermore, these bounds are shown to be tight in the case where the majority vote classifier makes its error mostly on low margin regions [25]. [24] demonstrated that this technique outperforms conventional fixed-threshold pseudo-labeling strategies on different multi-class classification problems.
2.3.5 Adaptive Thresholding Methods
@@
To address the limitations of confidence thresholding in self-training, [87] proposed an approach by analyzing the relationship between the ideal threshold and the model’s learning status. This approach adapts the confidence threshold self-adaptively based on the model’s progress. Additionally, it includes a self-adaptive class fairness regularization penalty to encourage diverse predictions during early training stages. In this line, [16] proposed a method to address the quantity-quality trade-off in pseudo-labeling. This approach aims to maintain both the quantity and quality of pseudo-labels during training, thereby effectively utilizing unlabeled data. This approach employs a truncated Gaussian function to weight samples based on their confidence, serving as a softer alternative to the confidence threshold. Furthermore, the authors introduced a uniform alignment approach to improve the utilization of weakly-learned classes.
Dynamic adjustment of thresholds reduces the risk of incorrect labels, and encourages diverse predictions during early training stages. However, it requires additional computational overhead to adapt thresholds and may introduce complexity in the training process.
2.4 Self-training with two classifiers
In co-training, examples are defined by two modalities that are comparable but not entirely correlated. Each view of an example is expected to contain complementary information about the data and if there are enough labeled training data, each of them is supposed to be sufficient for learning. The main principle is to learn a classifier on each view, taking initially the available labeled examples as the training set. Then, one of the classifiers assigns pseudo-labels to unlabeled data, and the other one uses them to retrain the model by including them into its training set. At each iteration, the classifiers alternately switch their roles, thereby co-training each other. As for self-training algorithms with a single classifier, this procedure continues until there are no more unlabeled instances to be pseudo-labeled. In practice, several studies artificially generated the two modalities for classification problems where examples are mono-viewed and described by a vector representation. These approaches create the two modalities out of one by selecting at random the set of features that should correspond to each modality from the initial set of features; and their efficiency was empirically proved on various applications [79].
Co-training and self-training share several fundamental principles. Both methods involve iteratively assigning pseudo-labels to unlabeled data and using these pseudo-labeled examples to improve the model. In self-training, a single model assigns pseudo-labels based on its confidence. In co-training, two or more models collaborate to assign pseudo-labels, with each model providing pseudo-labels for the data on which it is most confident. Furthermore, both approaches leverage unlabeled data to enhance the performance of the model(s) involved in training. In self-training, the model uses its own predictions to generate pseudo-labels for unlabeled data. In co-training, multiple models use their complementary strengths to generate pseudo-labels for unlabeled data, thereby improving the overall model performance by relying on confidence measures to select which unlabeled examples to pseudo-label. In self-training, the model’s confidence in its predictions determines which examples are pseudo-labeled. In co-training, each model’s confidence in its predictions determines which examples it will pseudo-label for the other model(s). Finally, both approaches involve iterative improvement of the model(s) through the incorporation of pseudo-labeled data. In self-training, the model is retrained with the augmented dataset that includes pseudo-labeled examples. In co-training, the models are retrained with the augmented dataset that includes pseudo-labeled examples provided by the other model(s).
Without splitting the input feature set, [17] proposed Cross Pseudo Supervision for semantic segmentation in images. This method employs two neural networks as supervised classifiers having the same images as inputs. Each neural-network is learned at every mini-batch by considering the pseudo-labels generated by the other network for unlabeled instances as ground-truths. In multi-task learning, [30] proposed to independently train specialized teachers using labeled datasets. These teachers then label an unlabeled dataset, creating a multitask pseudo-labeled dataset. Subsequently, a student model is trained using the pseudo-labeled dataset, employing multi-task learning to learn from various datasets and tasks simultaneously. Finally, the feature representations of the student model are evaluated across six vision tasks, including image recognition, to assess its perWe ance and generalization capabilities.
@@ 2.4 Self-training with two classifiers
The learnability of co-training was studied under the PAC framework [82], which also accounts for noise in the class labels of unlabeled examples caused by pseudo-labeling. The injection of noisy labels in the pseudo-labeling step is in fact inherent to any self-training algorithm. Taking into account noisy labels in training a model was first considered in supervised learning, following the paradigm of learning with an imperfect supervisor in which training data contains an unknown portion of imperfect labels [33]. Most of these studies tackle this problem from an algorithmic point of view, employing regularization or estimating mislabeling errors by modeling the transition probability between noisy and true labels [41, 60].
@@ 2.5 Self-training under Domain Shift
When intermediate domains are not given, it is important to ensure that pseudo-labeled target examples are reliable and are not biased towards the source data. While [100] approached this issue by carefully choosing a pseudo-labeling policy, [67] learn a representation via a tri-training scheme, in which the student is trained on target data pseudo-labeled by agreement of two teachers. [49] alternate between two gradient steps: (1) to train a source classification head that generates pseudo-labels, (2) to train a target classification head using pseudo-labeled data under the constraint that it predicts well on source data.
As the discrepancy between the source and the target can be large, the prediction confidence may exhibit a strong bias failing to distinguish between correct and wrong pseudo-labels. Therefore, several works focus specifically on model calibration and uncertainty estimation including the Monte-Carlo dropout [55], and prediction agreement of diversified linear heads [58].
@@ 3.1 Transductive learning
Transductive approaches also assume that the function class’s structure should reflect prior knowledge of the learning problem at hand, and that it should be built in such a way that the correct prediction of class labels of labeled and unlabeled training examples is contained in a function class $\mathcal{F}_{j}$ of small size with a high probability. In particular, the Transductive Support Vector Machines (TSVM) [36] developed for the binary case is based on this paradigm. The approach looks for the optimal hyperplane in a feature space that separates the best labeled examples while avoiding high density areas. TSVM does this by building a structure on a function class $\mathcal{F}$ and sorting the outputs of unlabeled samples by their margins. The solutions to the associated optimization problem are the pseudo-labels of unlabeled examples for which the resulting hyperplane separates the examples of both labeled and unlabeled training sets with the largest margin.
@@
Transductive learning is particularly useful in the context of proprietary and closed APIs in natural language processing, including few-shot classification. [20] proposed a scenario where pre-trained model embeddings are served through a gated API with compute-cost and data-privacy constraints, and introduces transductive inference as a solution. The authors present a new parameter-free transductive regularizer based on the Fisher-Rao loss, which fully utilizes unlabeled data without sharing labels with the API provider, and demonstrate its superiority through extensive experiments. More recently, [54] studied the problem of adversarially robust learning in the transductive setting. For classes with bounded VC dimension, the authors propose a simple transductive learner that correctly labels adversarially perturbed test examples with a robust error rate linear in the VC dimension, providing an exponential improvement over the best-known inductive setting bounds, albeit with a more restrictive notion of optimal robust error.